# Robust Kernel Hypothesis Testing under Data Corruption

### v1

@@ 5 Experiments
Figure 1: Two-sample experiments robust up of $r$ corrupted samples. A robust test needs to control the rejection rate by $\alpha=0.05$ when fewer than $r$ samples are corrupted. (Top row: Gaussian mean shift) Both samples are originally i.i.d. drawn from $\mathsf{Gaussian}(0,\nicefrac{{1}}{{10}},50)$, entries of one sample are corrupted being replaced by samples from $\mathsf{Gaussian}(1000,\nicefrac{{1}}{{10}},50)$. (Bottom row: IMDb movie reviews) Both samples originally consist of movie reviews (using a bag of 3330 words representation). Corrupted entries for one sample are replaced by samples from $\mathsf{Geometric}(0.05,3330)$.
@@
In this section, we provide simulation results for two-sample and independence testing on both synthetic and real-world data, focusing on the robustness properties of the designed tests. Our initial focus is to demonstrate that the robust tests control the type I error whenever $r$ or fewer entries have been corrupted, a task which the non-robust tests fail to accomplish. Assuming this control, the aim is then to have the test power, e.g., to detect when strictly more than $r$ samples have been corrupted.
@@
We stress that our methods are calibrated to control the type I error even under the most severe corruption scenarios. Under the strongest $c$ corruption scenario, we expect the rejection rate of our robust tests with robustness parameter $r$ to be exactly at $\alpha$ when $c=r$ (as the bounds in Equations 2 and 3 are tight in this setting). We also expect the test power to start increasing as soon as $c>r$. In practice, and in our experiments, however, we typically encounter milder corruption scenarios under which our robust tests are conservative for $c\leq r$, resulting in the type I error substantially smaller than $\alpha$. This implies a delay in the observed increase of test power, which sometimes happens only for values of $c$ much larger than $r$.
@@
In the two-sample setting, we compare the rejection rates achieved by the robust dcMMD and dpMMD tests, as well as by the non-robust permutation-based MMD test. Similarly, we compare the performance of all three dcHSIC, dpHSIC and HSIC tests for the independence testing problem. All tests are run using Gaussian kernels and 500 permutations. For both frameworks, we consider synthetic Gaussian simulations (mean shift and mixture), and experiments based on the real-world IMDb ‘Large Movie Review Dataset’ of [44] (see captions of Figures 1 and 2 for experimental details, rejection rates plotted are averaged over 200 repetitions).
@@
The same trends are observed across all experiments of Figures 1 and 2. First, we note that all tests control the type I error by $\alpha=0.05$ in the uncorrupted setting. As soon as some samples are corrupted, the MMD and HSIC tests achieve rejection rate 1, i.e., these tests are not robust to data corruption. In contrast, we observe that the DC and DP tests control the rejection rate to be lower than $\alpha$ when at most $r$ samples are corrupted, as theoretically guaranteed by Lemmas 2 and 1.
@@
In the alternative regime with more than $r$ corrupted samples, the DC and DP tests behave differently, with DC clearly always achieving higher power. The DC rejection rate rapidly increases to 1 after some threshold is reached in the number of corrupted samples, which is a desired property: the rejection rate is controlled to be lower than $\alpha$ when robustness is needed, and the power becomes 1 when robustness is no longer required. We observe that the DP rejection rate starts increasing around the same threshold as for DC, but, unlike DC, its growth is slow and gradual. We regard this as an empirical price to pay for having to satisfy the additional differential privacy constraint.
@@
As aforementioned, the fact that this threshold can sometimes be larger than $r$ can be explained by the fact that the tests are designed to be robust to any $r$ corruption processes, including stronger ones than those used in the experiments. Nonetheless, we highlight the strong performance of the DC tests. For example, dcMMD is able to detect differences in distributions using only $2000$ samples while being robust to the corruption of up $800$ of these samples (see Figure 1), which is remarkable.
@@
Figure 2: Independence experiments robust up of $r$ corrupted samples. A robust test needs to control the rejection rate by $\alpha=0.05$ when fewer than $r$ samples are corrupted. (Top row: Gaussian mixture) Paired samples $(X,Y)$ are originally i.i.d. drawn from two $\mathsf{Gaussian}(0,\nicefrac{{1}}{{10}},50)$. Corrupted samples are replaced by $(X,X+\epsilon)$ where $\epsilon\sim\mathsf{Gaussian}(0,\nicefrac{{1}}{{10}},50)$ and where $X\sim\mathsf{Gaussian}(s1000,\nicefrac{{1}}{{10}},50)$ with $s=1$ for half of the corrupted samples and $s=-1$ for the other half. (Bottom row: IMDb movie reviews) Paired samples $(X,Y)$ originally consist of two independent reviews (represented using a bag of 3330 words). Corrupted samples are replaced by $(X+s,X+s+\epsilon)$ where $X\sim\mathsf{Geometric}(0.05,3330)$, $\epsilon\sim\mathsf{Gaussian}(0,\nicefrac{{1}}{{10}},3330)$ and with $s=0$ for half of the corrupted samples and $s=5$ for the other half.
@@ 6 Conclusion
In this paper, we introduced two general approaches for constructing robust permutation tests under data corruption. The first method leverages recent advances in differentially private tests, achieving both robustness and privacy guarantees simultaneously. The second method avoids the use of random noise injection, making it more suitable for settings where reproducibility is critical. By employing the permutation principle, under the assumption of exchangeability, both approaches ensure non-asymptotic validity under data corruption. In terms of power guarantees, we established the minimal conditions under which the proposed tests maintain consistency under data corruption against any fixed alternatives. We illustrated our general frameworks in the context of kernel two-sample and independence testing, and showed that our kernel robust tests achieve minimax optimal power in terms of the MMD and HSIC metrics. Finally, empirical results were presented to illustrate the finite-sample performance of our methods.
@@
Limitations and future work. Our work opens up several fruitful directions for future research. As illustrated in our empirical studies, the proposed methods might be conservative in certain settings, as they are designed to safeguard testing errors against worst-case attack scenarios. It would be interesting to consider milder and potentially more structured attack scenarios (e.g., Huber’s $\epsilon$-contamination model), and to refine our algorithms to mitigate this conservative nature. Furthermore, expanding our methods to other types of test statistics and different testing problems is certainly worth exploring. One can also attempt to develop computationally efficient robust tests by exploiting recent advancements in time-efficient kernel tests [e.g., 51, 21]. Lastly, studying minimax rates for robust testing in terms of other metrics, such as the Wasserstein and $L_{p}$ metrics, presents an interesting avenue for future work.
Appendix A Experiments resources
@@
Reproducible anonymised code in JAX [9] is available at github.com/antoninschrab/dckernel under the MIT License. Experiments were run on a 24Gb RTX A5000 GPU, total compute time is of the order of ten hours. For the DP tests, we adapt the implementation of [38] published under the MIT License. We use the publicly available IMDb dataset of [44].
@@ Appendix B Assumptions
    Assume that the kernels $k$ and $\ell$ are characteristic, non-negative, and bounded everywhere by constants $K$ and $L$ respectively, i.e., $0\leq k(y,y^{\prime})\leq K$ for all $y,y^{\prime}\in\mathbb{S}_{Y}$ and $0\leq\ell(z,z^{\prime})\leq L$ for all $z,z^{\prime}\in\mathbb{S}_{Z}$.
@@
As discussed in [38], commonly used kernels, such as the Gaussian kernel and the Laplace kernel, satisfy the above assumptions. We next describe the assumption for $B$, the number of permutations, used in Theorems 1 and 2.
@@
We make the following assumptions on the number of permutations $B$ for Theorem 1 and Theorem 2:
@@
The reason for the difference in the required number of permutations between the DP test and the DC test stems from their reliance on different techniques. Specifically, we can guarantee that $M_{0},M_{1},\ldots,M_{B}$ in Algorithm 1 are all distinct with probability one, due to the injection of continuous noise to the test statistics. This property allows us to employ the multiplicative Chernoff inequality as in [38, Lemma 21]. On the other hand, there is no guarantee that $T_{0},T_{1},\ldots,T_{B}$ are distinct in Algorithm 2, for which we apply the Dvoretzky–Kiefer–Wolfowitz inequality to analyse the random permutation distribution, as in [50, Proposition 4]. If we randomly break ties in $T_{0},T_{1},\ldots,T_{B}$, the condition for dcMMD/dcHSIC can be improved to $6\alpha^{-1}\log(2/\beta)$.
Appendix C Proofs
@@
Let $X\sim P$ and $Y\sim Q$. We use $d_{\mathrm{TV}}(P,Q)$ to denote the total variation (TV) distance between $P$ and $Q$. With an abuse of notation, we also use $d_{\mathrm{TV}}(X,Y)$ to denote $d_{\mathrm{TV}}(P,Q)$. As in the main text, we add the subscript $P$ in $\mathbb{P}_{P}$ to emphasise that the underlying data are generated from the distribution $P$. We often omit the dependence on $P$ in $\mathbb{P}_{P}$ whenever it is implicitly clear from the context.
@@ C.1 Proof of Lemma 1
(i) The non-asympotic validity of the adjusted DP test in the data corruption framework follows from Equation 2 since the test using the uncorrupted data is non-asymptotically valid (in the usual framework) and differentially private [38, Theorems 1 and 2].
(ii) By the DP group privacy property [24, Theorem 2.2], we get the type II error bounded above as
$\mathbb{P}_{P_{1}}\bigl{(}\textrm{DP fails to reject $\mathcal{H}_{0}$}\,\,|\,\,\textrm{corrupted data}\bigr{)}$ $\leq\leavevmode\nobreak\ e^{r\varepsilon}\mathbb{P}_{P_{1}}\bigl{(}\textrm{DP fails to reject $\mathcal{H}_{0}$}\,\,|\,\,\textrm{uncorrupted data}\bigr{)}$
$\leq\leavevmode\nobreak\ e^{\nu}\mathbb{P}_{P_{1}}\bigl{(}\textrm{DP fails to reject $\mathcal{H}_{0}$}\,\,|\,\,\textrm{uncorrupted data}\bigr{)}$
By [38, Theorem 3], the above bound converges to zero as $n\rightarrow\infty$ under the conditions of Theorem 1. This proves that the DP test is consistent in power. ∎

### v3

@@ 6 EXPERIMENTS
Figure 1: Two-sample experiments robust up to $r$ corrupted samples. To have valid level, a robust test needs to control the rejection rate by $\alpha=0.05$ when fewer than $r$ samples are corrupted. To be powerful, the robust test needs to have a high rejection rate when more than $r$ samples are corrupted. (Top row: Gaussian mean shift) Both samples are originally i.i.d. drawn from $\mathsf{Gaussian}(0,\nicefrac{{1}}{{10}},50)$, entries of one sample are corrupted being replaced by samples from $\mathsf{Gaussian}(1000,\nicefrac{{1}}{{10}},50)$. (Bottom row: IMDb movie reviews) Both samples originally consist of movie reviews (using a bag of 3330 words representation). Corrupted entries for one sample are replaced by samples from $\mathsf{Geometric}(0.05,3330)$.
@@
In this section, we provide simulation results for two-sample and independence testing on both synthetic and real-world data, focusing on the robustness properties of the designed tests. Our initial focus is to demonstrate that the robust tests control the type I error whenever $r$ or fewer entries have been corrupted, a task which the non-robust tests fail to accomplish. Assuming this control, the aim is then to have test power, e.g., to detect when strictly more than $r$ samples have been corrupted.
@@
We stress that the DC and DP methods (Algorithms 1 and 2) are calibrated to control the type I error even under the most severe corruption scenarios. Under the strongest $c$-corruption scenario where $c$ denotes the number of corrupted data points, we expect the rejection rate of the robust tests with robustness parameter $r$ to be exactly at $\alpha$ when $c=r$ (as the bounds in Equations 1 and 2 are tight in this setting). We also expect test power to start increasing as soon as $c>r$. In practice, and in our experiments, however, we typically encounter milder corruption scenarios under which the robust tests are conservative for $c\leq r$, resulting in a type I error substantially smaller than $\alpha$. This implies a delay in the observed increase of test power, which sometimes happens only for values of $c$ much larger than $r$.
@@
In the two-sample setting, we compare the rejection rates achieved by our proposed robust dcMMD test against the dpMMD test adapted from Kim and Schrab, (2023), as well as by the non-robust permutation-based MMD test. Similarly, we compare the performance of all three dcHSIC, dpHSIC and HSIC tests for the independence testing problem. All tests are run using Gaussian kernels and 500 permutations. For both frameworks, we consider synthetic Gaussian simulations (mean shift and mixture), and experiments based on the real-world IMDb ‘Large Movie Review Dataset’ of Maas et al., (2011) (see captions of Figures 1 and 2 for experimental details, rejection rates plotted are averaged over 200 repetitions).
@@
The same trends are observed across all experiments of Figures 1 and 2. First, we note that all tests control the type I error at $\alpha=0.05$ in the uncorrupted setting. As soon as some samples are corrupted, the MMD and HSIC tests achieve rejection rate 1, i.e., these tests are not robust to data corruption. In contrast, we observe that the DC and DP tests control the rejection rate to be lower than $\alpha$ when at most $r$ samples are corrupted, as theoretically guaranteed by Lemmas 1 and 4.
@@
In the alternative regime with more than $r$ corrupted samples, the DC and DP tests behave differently, with our proposed DC tests clearly always achieving higher power than the DP tests adapted from Kim and Schrab, (2023). The DC rejection rate rapidly increases to 1 after some threshold is reached in the number of corrupted samples, which is a desired property: the rejection rate is controlled to be lower than $\alpha$ when robustness is needed, and the power becomes 1 when robustness is no longer required. We observe that the DP rejection rate starts increasing around the same threshold as for DC, but, unlike DC, its growth is slow and gradual. Overall, while both procedures control the type I error under data corruption, the DC tests are much more powerful than the DP tests.
@@
As aforementioned, the fact that this threshold can sometimes be larger than $r$ can be explained by the design of the tests to be robust to any $r$ corruption processes, including stronger ones than those used in these experiments. Nonetheless, we highlight the strong performance of the DC tests. For example, dcMMD is able to detect differences in distributions using only $2000$ samples while being robust to the corruption of up to $800$ of these samples (see Figure 1), which is remarkable.
We note that the experiments in Figures 1 and 2 are presented as the null hypothesis holding and being corrupted towards some alternative hypothesis (e.g., reading the plots from left to right). They can equivalently be interpreted as some alternative hypothesis holding and being corrupted towards the null hypothesis (e.g., reading the plots from right to left).
@@
Figure 2: Independence experiments robust up to $r$ corrupted samples. To have valid level, a robust test needs to control the rejection rate by $\alpha=0.05$ when fewer than $r$ samples are corrupted. To be powerful, the robust test needs to have a high rejection rate when more than $r$ samples are corrupted. (Top row: Gaussian mixture) Paired samples $(X,Y)$ are originally i.i.d. drawn from two $\mathsf{Gaussian}(0,\nicefrac{{1}}{{10}},50)$. Corrupted samples are replaced by $(X,X+\epsilon)$ where $\epsilon\sim\mathsf{Gaussian}(0,\nicefrac{{1}}{{10}},50)$ and where $X\sim\mathsf{Gaussian}(s1000,\nicefrac{{1}}{{10}},50)$ with $s=1$ for half of the corrupted samples and $s=-1$ for the other half. (Bottom row: IMDb movie reviews) Paired samples $(X,Y)$ originally consist of two independent reviews (represented using a bag of 3330 words). Corrupted samples are replaced by $(X+s,X+s+\epsilon)$ where $X\sim\mathsf{Geometric}(0.05,3330)$, $\epsilon\sim\mathsf{Gaussian}(0,\nicefrac{{1}}{{10}},3330)$ and with $s=0$ for half of the corrupted samples and $s=5$ for the other half.
@@ 7 CONCLUSION
In this paper, we introduced a general approach for constructing robust permutation tests under data corruption. This method, called DC procedure and presented in Algorithm 1, avoids the use of random noise injection (required for the DP procedure of Kim and Schrab, (2023) presented in Algorithm 2), making it more suitable for settings where reproducibility is critical. By employing the permutation principle, under the exchangeability assumption, both approaches ensure non-asymptotic validity under $r$ data corruption.
@@
In terms of power guarantees, we established the minimal conditions under which the DC test maintains consistency under data corruption against any fixed alternatives. We illustrated our general frameworks in the context of kernel two-sample and independence testing, and showed that our kernel robust DC tests achieve minimax optimal power in terms of the MMD and HSIC metrics. Additionally, we demonstrated that in low corruption regimes, the robustness property can be attained without compromising the minimum separation rate in terms of power. Empirical results were presented to illustrate the finite-sample performance of our robust DC tests which significantly outperform the DP tests in terms of test power.
@@
Limitations and future work. Our work opens up several fruitful directions for future research. As illustrated in our empirical studies, the proposed methods might be conservative in certain settings, as they are designed to safeguard testing errors against worst-case attack scenarios. It would be interesting to consider milder and potentially more structured attack scenarios (e.g., Huber’s $\epsilon$-contamination model), and to refine our DC procedure to mitigate this conservative nature. Furthermore, expanding our methods to other types of test statistics and different testing problems is worth exploring. One can also attempt to develop computationally efficient robust tests by exploiting recent advancements in time-efficient kernel tests (e.g., Domingo-Enrich et al.,, 2023; Schrab et al., 2022b, ). Lastly, studying minimax rates for robust testing in terms of other metrics (e.g., Wasserstein and $L_{p}$ metrics) presents an interesting avenue for future work.
Acknowledgments
Antonin Schrab acknowledges support from the U.K. Research and Innovation (grant EP/S021566/1). Ilmun Kim acknowledges support from the Basic Science Research Program through the National Research Foundation of Korea (2022R1A4A1033384), and the Korea government (MSIT) (RS-2023-00211073).
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Yes
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Yes
    2. (b)
      Complete proofs of all theoretical results. Yes
    3. (c)
      Clear explanations of any assumptions. Yes
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Not Applicable (outputs are binary so error bars are deterministic given the number of repetitions)
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Yes
    2. (b)
      The license information of the assets, if applicable. Yes
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. Yes
    4. (d)
      Information about consent from data providers/curators. Yes
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
Supplementary Materials
Robust Kernel Hypothesis Testing under Data Corruption
In this supplementary material, we discuss computational resources (Appendix A), present assumptions (Appendix B), derive proofs of the theoretical results presented in the main text (Appendix C), and offer a power analysis of the DP tests (Appendix D).
Appendix A EXPERIMENTAL RESOURCES
@@
Reproducible code in JAX (Bradbury et al.,, 2018) is available at https://github.com/antoninschrab/dckernel-paper under the MIT License. Experiments were run on a 24Gb RTX A5000 GPU, total compute time is of the order of ten hours. For the DP tests, we adapt the implementation of Kim and Schrab, (2023) published under the MIT License. We use the publicly available IMDb dataset of Maas et al., (2011).
@@ Appendix B ASSUMPTIONS
    Assume that the kernels $k$ and $\ell$ are characteristic, non-negative, and bounded everywhere by constants $K$ and $L$, respectively, i.e., $0\leq k(y,y^{\prime})\leq K$ for all $y,y^{\prime}\in\mathbb{S}_{Y}$ and $0\leq\ell(z,z^{\prime})\leq L$ for all $z,z^{\prime}\in\mathbb{S}_{Z}$.
@@
As discussed in Kim and Schrab, (2023), commonly used kernels, such as the Gaussian kernel and the Laplace kernel, satisfy the above assumptions. We next describe the assumption for $B$, the number of permutations, used in Theorems 1, 3, 2 and 4.
@@
We make the following assumptions on the number of permutations $B$ for Theorems 1, 3, 2 and 4:
@@
The reason for the difference in the required number of permutations between the DP test and the DC test stems from their reliance on different techniques. Specifically, we can guarantee that $M_{0},M_{1},\ldots,M_{B}$ in Algorithm 2 are all distinct with probability one, due to the injection of continuous noise to the test statistics. This property allows us to employ the multiplicative Chernoff inequality as in Kim and Schrab, (2023, Lemma 21). On the other hand, there is no guarantee that $T_{0},T_{1},\ldots,T_{B}$ are distinct in Algorithm 1, for which we apply the Dvoretzky–Kiefer–Wolfowitz inequality to analyse the random permutation distribution, as in Schrab et al., (2023, Proposition 4). If we randomly break ties in $T_{0},T_{1},\ldots,T_{B}$, the condition for dcMMD/dcHSIC can be improved to $6\alpha^{-1}\log(2/\beta)$.
Appendix C PROOFS OF DC PROCEDURE MINIMAX OPTIMALITY
@@
Let $X\sim P$ and $Y\sim Q$. We use $d_{\mathrm{TV}}(P,Q)$ to denote the total variation (TV) distance between $P$ and $Q$. With an abuse of notation, we also use $d_{\mathrm{TV}}(X,Y)$ to denote $d_{\mathrm{TV}}(P,Q)$. As in the main text, the subscript $P$ in $\mathbb{P}_{P}$ is used to emphasise that the underlying data are generated from the distribution $P$. We often omit the dependence on $P$ in $\mathbb{P}_{P}$ whenever it is implicitly clear from the context. We stress that the probabilities are also taken with respect to $\bm{\pi}$, that is, with respect to the uniformly random draw of permutations.
@@ C.1 Proof of Lemma 1
(i) To simplify the notation, define $U_{0}=T({\widetilde{\mathcal{X}}_{n}})$, $U_{i}=T({\widetilde{\mathcal{X}}_{n}}^{\bm{\pi}_{i}})$ for $i\in[B]$, (i.e., the test statistics based on the uncorrupted dataset ${\widetilde{\mathcal{X}}_{n}}$ and its permuted counterparts ${\widetilde{\mathcal{X}}_{n}}^{\bm{\pi}_{i}}$) and denote the $(1-\alpha)$-quantile of $U_{0},U_{1},\ldots,U_{B}$ as $\widetilde{q}$. Equipped with this notation, observe that $|U_{i}-T_{i}|\leq r\Delta_{T}$ for each $i\in[B]_{0}$, which follows by the repeated use of the triangle inequality and by using the definition of global sensitivity, i.e., $\Delta_{T}=\sup_{\bm{\pi}\in\bm{\Pi}_{n}}\sup_{\begin{subarray}{c}\mathcal{X}_{n},{\mathcal{Y}}_{n}\,:\,d_{\mathrm{ham}}(\mathcal{X}_{n},{\mathcal{Y}}_{n})\leq 1\end{subarray}}|T(\mathcal{X}_{n}^{\bm{\pi}})-T({\mathcal{Y}}_{n}^{\bm{\pi}})|$. This inequality yields
Since $U_{0},U_{1},\ldots,U_{B}$ are exchangeable under the null, we have $\mathbb{P}_{P_{0}}(T_{0}>q+2r\Delta_{T})\leq\mathbb{P}_{P}(U_{0}>\tilde{q})\leq\alpha$ for any $P_{0}\in\mathcal{P}_{0}$ (Romano and Wolf,, 2005, Lemma 1), which proves the validity result.
(ii) Using the same notation, we first note that the type II error of the DC test is bounded above as follows
where we use the fact that $|U_{i}-T_{i}|\leq r\Delta_{T}$ for each $i\in[B]_{0}$. Therefore, in order to prove that the DC test is consistent, it suffices to show that the above upper bound for the type II error converges to zero as $n\rightarrow\infty$ under the given conditions. Now, denoting $\overline{U}_{i}=U_{i}+4r\Delta_{T}$ for $i\in[B]_{0}$, by definitions of quantiles, we obtain
$=\mathds{1}(U_{0}>\widetilde{q}+4r\Delta_{T}).$
Using this alternative representation of the test in conjunction with Kim and Schrab, (2023, Lemma 8), it follows that